### v1

# Krylov complexity of thermal state in early universe

Abstract
In our work, we perform a detailed study of the Krylov complexity of the thermal state across the entire early universe, encompassing the inflation, radiation-dominated period, and matter-dominated period, which is for the single field inflation. We utilize both the closed system’s method and open system’s method to achieve this goal. To accurately calculate the Krylov complexity, we purified the thermal state, resulting in a pure state with two modes. Our analysis with both methods indicates that the Krylov complexity will increase during inflation, but will saturate at constant values during the radiation-dominated and matter-dominated periods, where the generation of particles via preheating leads to this evolution. Furthermore, our findings reveal that inflation behaves as a strong dissipative system, while the radiation-dominated and matter-dominated periods act as weak dissipative systems. The chaotic feature during these periods follows a similar trend to the Krylov complexity. This research has the potential to provide new insights into the exploration of Krylov complexity in cosmology.
1 Introduction
The concept of complexity arose from the holographic framework Stanford:2014jda , where they stated that the conformal field reaches the thermal equilibrium within a short temporal scale. Still, its corresponding bulk gravity arrives at a certain state with a much longer time. To solve this issue, Refs. Hartman:2013qma ; Liu:2013iza introduced the concept of computational complexity under the holographic framework, which responses to the evolution of spacetime, and the emergence of spacetime comes via quantum entanglement.
The very significant development of complexity in high-energy physics allows exploration without the holographic principle. First, Ref. Aaronson:2016vto defined computational complexity as the minimal number of logical procedures for a given task. As for the method of calculating the complexity, Nielson, $\it e.t.c$ pioneering the so-called the geometrical method for calculating the so-called circuit complexity Nielsen:2005mkt ; Nielsen:2006cea ; Dowling:2006tnk . The second method is explicitly related to the ”Fubini study” distance of the geometry of information Chapman:2017rqy . Physically speaking, these two methods are utilized for the circuit complexity, depending on the parametric manifold and the relevant metric for the first and second method, respectively. The third method of calculating the complexity is called the Krylov complexity, which is free of ambiguities compared with the previous two. Due to the uniqueness of Krylov complexity, one can compare the evolution of Krylov complexity with the circuit complexity to obtain the differences and similarities. Based on the Krylov complexity, Ref. Parker:2018yvk proposed a universal operator growth. Meanwhile, one can investigate the Krylov complexity, Krylov entropy, and its corresponding Lanczos coefficient in light of the Lanczos algorithm when the quantum state is given viswanath1994recursio . Ref Aguilar-Gutierrez:2023nyk has shown the relation between the circuit complexity and Krylov complexity, which are not compatible with each other. But under certain conditions, Krylov complexity can be proportional to circuit complexity (Fubini distance) Caputa:2021sib .
In recent years, the Krylov complexity has made rapid progress in condensed matter and high-energy physics. Ref. Muck:2022xfc has systematically investigated the Krylov complexity in various polynomials dubbed as the inner product. The Krylov complexity has been applied to the SKY model Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk , the generalised coherent state Patramanis:2021lkx , Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , conformal field theory Dymarsky:2021bjq ; Caputa:2021ori , topological phases of matter Caputa:2022eye . The chaos can be understood as a delocalization in the Krylov space Dymarsky:2019elm . Thereafter, Ref. Bhattacharjee:2022vlt has shown that the exponential growth of Krylov complexity can be observed in integrable systems with saddle-dominated scrambling, where Ref. Huh:2023jxt verified all the universal features of spread complexity in this framework. The very recent developments of Krylov complexity can be found in Refs. Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj . Especially, the detailed implication of Krylov complexity in various fields can be found in Ref. Nandy:2024htc .
This paper aims to investigate Krylov complexity in the early universe, encompassing the inflationary period, matter-dominated period (MD), and radiation-dominated period (RD). There has been research in cosmology addressing complexity in this context Choudhury:2020hil ; Bhargava:2020fhl ; Lehners:2020pem ; Bhattacharyya:2020rpy ; Adhikari:2021ked . The research has shown oscillations during inflation, and the evolution trend of complexity thereafter is similar post-inflation Liu:2021nzx ; Li:2021kfq ; Li:2023ekd . When incorporating the effects of various quantum gravitational frameworks, complexity demonstrates irregular oscillation during inflation Li:2023ekd . However, most research has focused on circuit complexity, with only a few studies on early universe Krylov complexity. Ref. Adhikari:2022oxr investigated Krylov complexity with various values of the speed of sound in inflation using a closed system approach. However, our universe is an entirely open system. Based on the method of an open system, Ref. Li:2024kfm examined the Krylov complexity with the modified dispersion relation in inflation, finding that the inflationary period behaves as a strong dissipative system and the overall trend of Krylov complexity is increasing. The key aspect here is constructing the wave function using the approach of open system, as presented by Ref. Bhattacharya:2022gbz , which also serves as our starting point for constructing the wave function for the thermal state.
In the early universe, the temperature is very high which lead to the expansion of the unverse and the generation of particles. Consequently, it is more dependable and practical to consider thermal effects for the wave function. Ref. Kar:2021nbm has defined a thermal Krylov complexity. In the cosmological field, Ref. Wang:2022mix has investigated the circuit complexity in the early universe, which indicated that the circuit complexity will decrease in the early universe. Based on it, we investigated the inflationary Krylov complexity of thermal state Li:2024iji . It should be noticed that the Krylov complexity can be extended into the quantum field theory (scalar field) with arbitrary temperature He:2024xjp . In light of Li:2024iji , we will continue to investigate the Krylov complexity in the whole early stage of the universe for the standard single field inflation.
This paper will be organized as follows. In Sec. 2, we will introduce the Lanczos algorithm in an approach of closed system. Sec. 3 will give a brief introduction to the early universe, including the inflation, RD, and MD, and the formula of scale factor in these three periods, respectively. In Sec. 4 will give the thermal state. Sec. 6 will investigate the Krylov complexity in a method of closed system. Sec. 7 will give the Lanczos coefficient and Krylov complexity with the approach of open system. Finally, the conclusion and outlook will be given in Sec. 8. In this paper, we will unify the units in the Planck units.
2 Lanczos algorithm with an approach of closed system
In this section, we will introduce the Lanczos algorithm with an approach of closed system. First, we need to define the operator $\mathcal{O}(t)$ in Heisenberg’s picture as follows,
$\begin{split}\partial_{t}\mathcal{O}(t)=i[H,\mathcal{O}(t)]\end{split},$ (1)
the corresponding solution is
$\begin{split}\mathcal{O}(t)=e^{iHt}\mathcal{O}e^{-iHt}.\end{split}$ (2)
And it could define the Liouvillian super-operator $\mathcal{L}_{X}$ as $\mathcal{L}_{X}Y=[X,Y]$ (the same commutator with (1)). Under the Liouvillian super-operator, one can obtain $\mathcal{O}(t)$ as follows,
$\begin{split}\mathcal{O}(t)=e^{i\mathcal{L}t}\mathcal{O}=\sum_{n=0}^{\infty}\frac{(it)^{n}}{n!}\mathcal{L}^{n}\mathcal{O}(0)=\sum_{n=0}^{\infty}\frac{(it)^{n}}{n!}\tilde{\mathcal{O}}_{n}.\end{split}$ (3)
where $\mathcal{O}(t)$ can be considered as the wave function, utilizing the $\mathcal{L}^{n}\mathcal{O}=\tilde{\mathcal{O}}_{n}=[H,\tilde{\mathcal{O}}_{n-1}]$ to form the Hilbert space,
$\begin{split}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}),\mathcal{L}^{1}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{1}),\mathcal{L}^{2}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{2}),\mathcal{L}^{3}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{3})...\end{split}$ (4)
Noticing that the basis is not orthogonal. Therefore, we could construct the orthogonal basis making using of the Lanczos algorithm, dubbed as the Krylov basis, where the first two terms are
$\begin{split}\mathcal{O}_{0}=|{\tilde{\mathcal{O}}}_{0})=\mathcal{O},|\mathcal{O}_{1})=b_{1}^{-1}\mathcal{L}|\mathcal{O}_{0}),\end{split}$ (5)
where $b_{1}=\sqrt{(\tilde{\mathcal{O}}_{0}\mathcal{L}|\mathcal{L}\tilde{\mathcal{O}}_{0})}$ is the normalized vector, by the iterative relation, another orthogonal basis can be formed,
$\begin{split}|A_{n})=\mathcal{L}|\mathcal{O}_{n-1})-b_{n-1}|\mathcal{O}_{n-2})\end{split}$ (6)
with
$\begin{split}|\mathcal{O}_{n})=b_{n}^{-1}|A_{n}),b_{n}=\sqrt{(A_{n}|A_{n})}\end{split}$ (7)
As $b_{n}=0$, this iterative relation (6) will stop and will generate the finite orthogonal Krylov basis, where $b_{n}$ is the so-called Lanczos coefficient. Thus, we could expand $\mathcal{O}$ as follows
$\begin{split}\mathcal{O}(t)=e^{i\mathcal{L}t}\mathcal{O}=\sum_{n=0}^{\infty}(i)^{n}\phi_{n}(t)|\mathcal{O}_{n}),\end{split}$ (8)
[…]
$\begin{split}\partial_{t}\phi_{n}(t)=b_{n}\phi_{n-1}-b_{n-1}\phi_{n+1},\end{split}$ (9)
According to this, the Krylov complexity can be obtained from the following formula
$\begin{split}K=\sum_{n}n\left|\phi_{n}\right|^{2}.\end{split}$ (10)
In addition, the Lanczos coefficient is linearly bounded via Parker:2018yvk ,
$b_{n}\leq\alpha n+\gamma$ (11)
where $\alpha$ and $\gamma$ that encode different information in different models, where it should be noted that the maximal value of $b_{n}$ denotes the maximal chaos of the dynamical system. In terms of this maximum chaos, the Lyapunov index $\gamma$ can be used to describe the chaotic growth of the system as follows,
$\gamma=2\alpha.$ (12)
Before investigating the Krylov complexity, we need the information of scale factor $a(\eta)$ in distinctive periods.
3 Some basics of early universe
In this section, we will give some fundamental knowledge for the early universe, including inflation, RD, and MD according to Baumann:2009ds , which is characterized by the scale factor $a(\eta)$ where $\eta$ is the conformal time.
3.1 Scale factor in various epoch
The starting point is the Friedman-Lemaitre-Robertson-Walker background metric, which is a good approximation for the evolution of the whole universe,
$ds^{2}=a(\eta)^{2}(-d\eta^{2}+d\vec{x}^{2}),$ (13)
where $a(\eta)$ is the so-called scale factor in terms of conformal time. Being armed with this metric, we could discuss its formula in various periods. Following Ref. Baumann:2009ds , we only consider a single component universe, which means that the whole universe experiences inflation, RD and MD, respectively. Parameterising this single component universe is the so-called equation of state $w_{I}=\frac{P_{I}}{\rho_{I}}$ where $I$ denotes these three distinctive periods. To obtain the explicit relation between the scale factor and the conformal time, we need the comoving distance which is defined as
$\chi_{ph}(\eta)=\int_{t_{i}}^{t}\frac{da}{a}=\int_{\ln{a_{i}}}^{\ln{a}}(aH)^{-1}d\ln{a},$ (14)
where $\chi_{ph}(\eta)$ is the comoving distance in the event horizon of particle and it can characterized by the so-called Hubble radius $(aH)^{-1}$. In light of $w_{I}$, one can rewrite the Hubble radius as $(aH)^{-1}=H_{0}^{-1}a^{\frac{1}{2}(1+3\omega)}$ where we use $w$ to take place of $w_{I}$ in the following investigations. Integrating both sides of comoving distance (14), we could get the following relation
$\eta\equiv\frac{2H_{0}^{-1}}{(1+3\omega)}a^{\frac{1}{2}(1+3\omega)}.$ (15)
Consequently, we can get the relation in various periods as following,
$\begin{split}\eta=\left\{\begin{matrix}&aH_{0}^{-1}(\omega=\frac{1}{3}),RD\\ &2\sqrt{a}H_{0}^{-1}(\omega=0),MD\end{matrix}\right.\end{split}$ (16)
As for the inflation, we already have $\eta=\frac{-1}{aH_{0}}$. Being armed with these three relations, we could investigate the Krylov complexity in the whole early universe.
3.2 Inflation, RD and MD
The whole early universe is composed of inflation, RD, and MD, respectively, which is characterized by the scale factor as mentioned in Sec. 3.1. Remember that there is an exponential expansion after the Big Bang (the creation of singularity), where the curvature perturbation will transit from the quantum level to the classical level after the horizon exits. After that, the universe will experience the so-called RD whose energy density is depicted by $\rho_{\rm rad}(t)=\rho_{\rm rad_{0}}\frac{a_{0}^{4}}{a(t)^{4}}$ with $\rho_{\rm rad_{0}}$ and $a_{0}$ are initial values of the energy density of RD and scale factor, respectively, where the particles in RD are relativistic. As the temperature decreases, the particles will become more and more un-relativistic, consequently, the universe will enter the so-called MD whose energy density is $\rho_{\rm mat}(t)=\rho_{\rm mat_{0}}\frac{a_{0}^{3}}{a(t)^{3}}$. Especially for the moment at $\rho_{\rm rad}(t)=\rho_{\rm mat}(t)$, it means that the universe will begin the MD epoch. During the whole section, we hope to help the reader to better understand the basic picture of the early universe, which is characterized by the scale factor.
4 The wave function of thermal state
Physically speaking, the Hamiltonian of single-field inflation can be considered as an inverted harmonic oscillator Li:2024kfm . Meanwhile, we will investigate the Krylov complexity for the thermal state.
$\begin{split}\hat{\rho}=\frac{1}{\sqrt{z}}\sum^{\infty}_{n=0}e^{-\beta E_{n}}|n\rangle\langle n|\end{split}$ (17)
In the context of bosonic oscillators, where $\beta=\frac{1}{k_{B}T}$ and $E_{n}=n\omega$ ($\omega$ is the frequency), we will consider the thermal state (17) as a pure state following the approach in Jefferson:2017sdb ; Ali:2018fcz , using the method described in Haque:2021kdm . Given that the thermal state is a mixed state, we need to create the purification of $\hat{\rho}$, which we refer to as a pure state $|\Psi\rangle$, in an enlarged Hilbert space $\mathcal{H}_{pure}=\mathcal{H}\otimes\mathcal{H}_{anc}$, where $\mathcal{H}_{anc}$ represents the ”ancillary” degrees of freedom. When we trace $|\Psi\rangle$ over $\mathcal{H}_{anc}$, we obtain the mixed state $\hat{\rho}_{mix}$, indicating that $|\Psi\rangle$ serves as the purification of the thermal state (17).
It is important to note that the observable is preserved even in the purification process, as evidenced by $\hat{\mathcal{O}}=\rm Tr_{anc}\langle\Psi|\mathcal{O}|\Psi\rangle=\rm Tr\big{(}\hat{\rho}_{mix}\hat{\mathcal{O}}\big{)}$, which maintains the reliability of calculating the Krylov complexity using this approach. However, it’s worth noting that the construction of $|\Psi\rangle$ is not unique due to the arbitrary choices of $\mathcal{H}_{anc}$. In this paper, we specifically duplicate the original Hilbert space, i.e., $\mathcal{H}_{anc}=\mathcal{H}$. With this treatment, we can obtain the classical thermofield double state (TFD).
$\left|TFD\right\rangle=\frac{1}{\sqrt{Z}}\sum^{\infty}_{n=0}e^{-\beta E_{n}/2}\left|n\right\rangle\otimes\left|n\right\rangle_{anc}$ (18)
Indeed, one can multiply a phase factor for TFD (18) which will not change its structure of wave function. Thus, one possible purification of the thermal state can be dented as follows,
$\left|\Psi\right\rangle_{\phi}=\left|TFD\right\rangle_{\phi}=\frac{1}{\sqrt{Z}}\sum^{\infty}_{n=0}(-1)^{n}e^{-2in\phi_{k}}e^{-\beta E_{n}/2}\left|n\right\rangle\otimes\left|n\right\rangle_{anc},$ (19)
where $(-1)^{n}e^{-2in\phi_{k}}$ is the phase factor, and it can identified as a two-mode squeezed vacuum state,
$\left|\Psi\right\rangle_{\phi}=\frac{1}{\cosh{r_{k}}}\sum^{\infty}_{n=0}(-1)^{n}e^{-2in\phi_{k}}\tanh^{n}{r_{k}}\left|n\right\rangle\otimes\left|n\right\rangle_{anc}=\hat{S}_{sq}(r_{k},\phi_{k})\left|0\right\rangle\otimes\left|0\right\rangle_{anc}.$ (20)
The relationship between the squeezing parameter $r_{k}$ and the oscillator frequency $(\omega)$ and temperature $(T)$ is as follows: $\beta\omega=-\ln{\tanh^{2}{r_{k}}}$, $\cosh r_{k}=\frac{1}{\sqrt{1-e^{-\frac{\omega}{k_{B}T}}}}$. So, we could obtain the wave function,
$\left|\Psi\right\rangle_{\phi}=\sqrt{1-e^{-\frac{\omega}{k_{B}T}}}\sum^{\infty}_{n=0}e^{-2in\phi_{k}}e^{-\frac{n\omega}{2k_{B}T}}\left|n\right\rangle\otimes\left|n\right\rangle_{anc}.$ (21)
This wave function is a pure state which will be used to investigate its corresponding Krylov complexity and Krylov entropy. In our case, we have $|\mathcal{O}_{n}\rangle=|n\rangle\otimes|n\rangle_{anc}$. Furthermore, it will also be implemented for constructing the wave function with the approach of open system based on Bhattacharya:2022gbz .
5 Evolution of temperature and squeezed angle
In this section, we will follow the traditional procedure to investigate the evolution of squeezed parameters, namely $r_{k}$ and $\phi_{k}$. In our case, the situation is different since $r_{k}$ is determined by temperature and frequency whose wave function is (21). In some sense, we transit the evolution of $r_{k}$ to the evolution of temperature since the factor of frequency could be absorbed into temperature due to $\beta\omega=-\ln\tanh^{2}r_{k}$ ($\beta=1/(k_{B}T)$).
Before the analysis of temperature and $\phi_{k}$, we will investigate the Krylov complexity of single-field inflation. Our starting metric is (13). Being different from the previous work Li:2024kfm , we will not adopt the Mukhanov-Sasaki variable since it contains the contribution of potential which cannot show the explicit impact on the Krylov complexity. Therefore, we will directly perturb the inflaton as follows,
$\phi(\eta,x)=\bar{\phi}(\eta)+\frac{v(\eta,x)}{a(\eta)},$ (22)
where $\bar{\phi}(\eta)$ is the background field of inflaton and $v(\eta,x)$ is the perturbation of inflaton. And our action of inflaton can read as
[…]
where $g^{\mu\nu}$ is the metric corresponding to (13). Plugging the perturbation of inflaton (22) into action (23) upon the second order, one can get the following action
[…]
where $V,_{\phi\phi}=\frac{d^{2}V}{d\phi^{2}}$ and we have performed a partial integration for the term $v^{2}$. The important point is that we can include the potential in this action, unlike in the action containing the Mukhanov-Sasaki variable, where the term $a^{2}V_{,\phi\phi}v^{2}$ can be ignored due to the slow-roll conditions. However, it should be taken into account in RD or MD due to the violation of slow-roll approximations. In general, the action given in equation (24) can be applied to the entire early universe, including inflation, RD, and MD. As for the potential of action (24), one could see that the formula of potential can be arbitrary which will lead to the impossibility of further investigations. Let us recall the process of RD and MD, the particle will be generated during these two processes that are dubbed as the preheating Kofman:1997yn ; Kofman:1994rk , where the energy of inflaton will be transferred into the standard model’s particles, such as the Higgs field, Fermions $\it e.t.c$. During this process, the potential will be oscillating around the minimum of the potential whose formula can be approximated to be
$V(\phi)=\frac{1}{2}m^{2}_{eff}\phi^{2},$ (25)
where $m_{eff}$ is the effective mass of inflaton and the background of $\phi$ (inflaton) will be transformed leading the varying of $\phi$. Being armed with this effective potential of inflaton, we could simulate the evolution of Krylov complexity.
Next, we need to transform action (24) into the Hamilton which will be denoted by the creation and annihilation operators. First, we need to define the conjugate momentum
$\pi(\eta,\vec{x})=\frac{\delta L}{\delta{v}^{\prime}(\eta,\vec{x})}={v}^{\prime}-\frac{{a}^{\prime}}{a}v.$ (26)
[…]
With the application of Fourier decomposition,
$\hat{v}(\eta,\vec{x})=\int\frac{d^{3}k}{(2\pi)^{3/2}}\sqrt{\frac{1}{2k}}(\hat{c}_{\vec{k}}^{\dagger}v_{\vec{k}}^{\ast}(\eta)e^{-i\vec{k\cdot}\vec{x}}+{\hat{c}_{\vec{k}}}v_{\vec{k}}e^{i\vec{k\cdot}\vec{x}}),$ (28)
$\hat{\pi}(\eta,\vec{x})=i\int\frac{d^{3}k}{(2\pi)^{3/2}}\sqrt{\frac{k}{2}}(\hat{c}_{\vec{k}}^{\dagger}u_{\vec{k}}^{\ast}(\eta)e^{-i\vec{k\cdot}\vec{x}}-\hat{c}_{\vec{k}}u_{\vec{k}}e^{i\vec{k\cdot}\vec{x}}).$ (29)
As a result, we could obtain the Hamiltonian in terms of creation and annihilation operators,
$\begin{split}\hat{H}_{k}&=(\frac{a^{2}}{2k}V_{,\phi\phi}+k)(\hat{c}_{anc}^{\dagger}\hat{c}_{anc}+\hat{c}_{\vec{k}}\hat{c}^{\dagger}_{\vec{k}})+(\frac{a^{2}}{2k}V_{,\phi\phi}+i\frac{{a}^{\prime}}{a})\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{anc}^{\dagger}\\ &+(\frac{a^{2}}{2k}V_{,\phi\phi}-i\frac{{a}^{\prime}}{a})\hat{c}_{\vec{k}}\hat{c}_{anc}\end{split}$ (30)
where $\hat{c}_{anc}$ corresponds to the $\hat{c}_{-\vec{k}}$ compared with our previous works Liu:2021nzx ; Li:2023ekd ; Li:2024kfm . One could discover that the the Hamilton (30) could nicely recover into the Hamilton for the inflation epoch,
$\begin{split}\hat{H}_{k}=k(\hat{c}_{anc}^{\dagger}\hat{c}_{anc}+\hat{c}_{\vec{k}}\hat{c}^{\dagger}_{\vec{k}})+i\frac{{a}^{\prime}}{a}\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{anc}^{\dagger}-i\frac{{a}^{\prime}}{a}\hat{c}_{\vec{k}}\hat{c}_{anc}\end{split}$ (31)
Our starting wave function is Eq. (21), and it satisfies with the Schrödinger equation,
$i\frac{d}{d\eta}\left|\Psi\right\rangle_{\phi}=\hat{H}_{k}\left|\Psi\right\rangle_{\phi},$ (32)
By using the Hamiltonian in Eq. (30), we can obtain the evolution equations of $T_{k}(\eta)$ and $\phi_{k}(\eta)$
$\ &{T}^{\prime}_{k}=-\frac{4k_{B}T_{k}^{2}}{\omega}\sinh{(\frac{\omega}{2k_{B}T_{k}})}[\frac{a{}^{\prime}}{a}\cos(2\phi_{k})+\frac{a^{2}}{2k}V_{,\phi\phi}\sin{(2\phi_{k})}]\\ &{\phi}^{\prime}_{k}=\frac{a^{2}}{2k}V_{,\phi\phi}+k+2\cosh{(\frac{\omega}{2k_{B}T_{k}})}[\frac{a{}^{\prime}}{a}\sin(2\phi_{k})-\frac{a^{2}}{2k}V_{,\phi\phi}\cos{(2\phi_{k})}].\end{split}$ (33)
Similarly, when we ignore the potential energy term, we can get the evolution equation in the inflation period.
$\ &{T}^{\prime}_{k}=-\frac{4k_{B}T_{k}^{2}}{\omega}\sinh{(\frac{\omega}{2k_{B}T_{k}})}\frac{a{}^{\prime}}{a}\cos(2\phi_{k})\\ &{\phi}^{\prime}_{k}=k+2\cosh{(\frac{\omega}{2k_{B}T_{k}})}\frac{a{}^{\prime}}{a}\sin(2\phi_{k}).\end{split}$ (34)
Being with the conformal time (16) in various epochs, we could explore the evolution of temperature $T_{k}$ and $\phi_{k}$ in terms of conformal time $\eta$.
Figure 1: The numerical solution of $T_{k}$ in terms of $\eta$ for three different periods (inflation, $RD$ and $MD$), in which we have set $k_{B}=1$, $\omega=1$, $H_{0}=10^{-2.5}$ and $m_{eff}=10^{-6}$ for simplicity. The momentum can be set various values in distinctive periods as showing in these panels.
Fig. 1 indicates the evolution of $T_{k}$ in the early universe, including inflation, RD, and MD. In this figure, we have set $\omega=1$ whose contribution can be absorbed into the temperature, $k_{B}=1$, $m_{eff}=10^{-6}$ and $H_{0}=10^{-2.5}$, where the values of $k$ are different in various periods as showing in distinctive panels. One could see that the temperature will increase, especially at the end of inflation. From the perspective of cosmology, we could conclude that more and more energy will be generated in inflation characterized by the temperature $T_{k}$. As for the RD and MD, the temperature will oscillate around some constant values since we already take account for the contribution of potential. Physically speaking, the particle will be generated during preheating (in MD) and is considered a highly non-equilibrium process. Although this process is non-equilibrium which is realized by the transformation of energy from inflaton to other particles, the energy of inflaton is also created via a vacuum that leads to the temperature approaching some constant values during RD and MD. Another interesting issue that should be noticed is the evolution of temperature concerning the momentum $k$. In some sense, the higher the temperature, the bigger the values of $k$. Our numeric shows the different evolution trends. In each panel of Fig. 1, one could see that the higher values of $k$, the lower values of temperature for each corresponding energy scale, as shown in the various values of $k$ in Fig. 1.
Figure 2: The numerical solution of squeezed angle $(\phi_{k})$ in terms of $\eta$ for three different periods(inflation, $RD$ and $MD$), in which we have set $k_{B}=1$, $\omega=1$, $H_{0}=10^{-2.5}$ and $m_{eff}=10^{-6}$ for simplicity. The momentum can be set various values in distinctive periods as showing in these panels.
In the later investigations, we will show that the $\phi(\eta)$ part will not contribute the Krylov complexity due to its definition. But for the completeness, we also give the numeric of $\phi_{k}$ in various periods, including inflation, RD and MD. Fig. 2 shows the numeric of $\phi_{k}$. The first trend is that the large values of $k$, the faster growth of $\phi_{k}$ in various periods which is different with evolution of $T_{k}$. In inflation, $\phi_{k}$ will grow at first and then it will approach to some constant values, thereafter $\phi_{k}$ will continue growing in RD and MD.
In Sec. 5, we simply introduce the perturbation of inflaton in single field inflation, which is not the Mukhanov-Sasaki variable. Once defining the perturbation of inflaton in Eq. (22), it could take account for the contribution of potential for inflaton. Following the standard procedure, we give the evolution equation of $T_{k}$ and $\phi_{k}$ as shown in Eq. (33). In the later investigations, we will show that the Krylov complexity is only determined by the temperature, where Fig. 1 clearly indicates the numeric of $T_{k}$. The temperature will grow in inflation, and it will approach some constant values of $T_{k}$ once taking the potential into account.
6 Krylov complexity with an approach of closed system
In this section, we will investigate the Krylov complexity and Krylov entropy (K-entropy) with the numeric of $T_{k}$ as shown in Fig. 1, where the calculation method is the closed system’s approach Liu:2021nzx ; Li:2021kfq ; Li:2023ekd .
6.1 Krylov complexity
In this section, we will use the wave function for exploring the Krylov complexity in closed system’s method. In light of (10), we could find the corresponding operator wave function as follows,
$\phi_{n}=\sqrt{1-e^{-\frac{\omega}{k_{B}T}}}(-1)^{n}e^{-2in\phi_{k}}e^{-\frac{n\omega}{2k_{B}T}}.$ (35)
Also according to the definition of Krylov complexity, we could explicitly obtain
$\begin{split}K&=\sum_{n=0}^{\infty}n|\phi_{n}|^{2}\\ &=\frac{e^{-\frac{\omega}{k_{B}T}}}{1-e^{-\frac{\omega}{k_{B}T}}}\\ &=\frac{1}{2}e^{-\frac{\omega}{2k_{B}T}}\sinh^{-1}{(\frac{\omega}{2k_{B}T})}.\end{split}$ (36)
Figure 3: The numeric of Krylov complexity in terms of $\eta$ for three different periods(inflation, $RD$ and $MD$),in which we have set $k_{B}=1$, $\omega=1$, $H_{0}=10^{-2.5}$ and $m_{eff}=10^{-6}$ for simplicity. The momentum can be set various values in distinctive periods as showing in these panels.
Based on the resulting formula of Krylov complexity, it is explicitly independent of $\phi_{k}$. According to this formula, we can obtain the numeric value of Krylov complexity as shown in Fig. 3. The trend is quite similar to the temperature as shown in Fig. 1, with only the corresponding amplitude being different. Throughout the entire inflationary process, the Krylov complexity is growing, which is consistent with our previous research Li:2024kfm . When taking the contribution of potential into account in RD and MD, the Kyrlov complexity will saturate at some values, oscillating around constant values, which is consistent with Barbon:2019wsy . Ref. Li:2024kfm has also claimed that the Krylov complexity will saturate at some constant values, as RD and MD are weak dissipative systems. In later investigations, we will confirm this point using the method of open systems.
6.2 Krylov Entropy
In a quantum system, entropy is defined to measure the disorder of a system. In the framework of Krylov complexity, the Krylov entropy (K-entropy) could test the level of disorder of curvature perturbation in our case. Similarly, with the definition of von Neumann entropy, we could also define the K-entropy following from Barbon:2019wsy ,
$S_{K}=-\sum^{\infty}_{n=0}|\phi_{n}|^{2}\ln|\phi_{n}|^{2},$ (37)
where $\phi_{n}$ is Eq. (35). After some algebra, we could obtain the resulting formula of K-entropy,
$\begin{split}S_{K}&=-\sum^{\infty}_{n=0}|\phi_{n}|^{2}\ln|\phi_{n}|^{2}\\ &=\frac{\omega}{2k_{B}T}e^{-\frac{\omega}{2k_{B}T}}\sinh^{-1}{(\frac{\omega}{2k_{B}T})}-\ln{[2\sinh{(\frac{\omega}{2k_{B}T})}]}+\frac{\omega}{2k_{B}T}.\end{split}$ (38)
Figure 4: The numeric of K-entropy in terms of $\eta$ for three different periods(inflation, $RD$ and $MD$),in which we have set $k_{B}=1$, $\omega=1$, $H_{0}=10^{-2.5}$ and $m_{eff}=10^{-6}$ for simplicity. The momentum can be set various values in distinctive periods.
Being armed with Eq. (38), we could numerically simulate the K-entropy by means of closed system’s method.
Fig. 4 clearly indicates the numeric of K-entropy whose trend is almost the same with Krylov complexity as shown in Fig. 3. Recalling the meaning of entropy, which is measuring the disorder of a chaotic system. The bigger the values of entropy, the more chaotic the system. Within this concept, we can see that inflation will be more and more chaotic. Once entering RD, the chaos will reach the maximal values. When more and more particles are generated via preheating, the chaos will evolve until another constant value.
In Sec. 6, we could see the entire evolution of Krylov complexity, where the Krylov complexity will increase in inflation and then it will saturate some values in RD and MD. As for the K-entropy, it represents the disorder of the universe, we could see that the chaos will be enhanced during inflation, and then the chaos will saturate some constant values, which means that the chaos will reach the maximum during preheating (in RD and MD).
7 Krylov complexity with the approach of open system
As we mentioned, the RD and MD are a highly non-equilibrium system due to the generation of particles. From another perspective, the universe is not invariant since reversing time leads to the non-conservation of energy. Thus, it is quite natural to generalize our method to the method of open system.
7.1 Lanczos coefficient and dissipative strength
In this section, we will follow the method of Bhattacharya:2022gbz to investigate which has been applied to our previous research Li:2024kfm . First, the general operator in Heisenberg’s picture can be written by,
$\begin{split}\mathcal{O}(\eta)=e^{i\mathcal{L}_{o}\eta},\end{split}$ (39)
where the Lindbladian $\mathcal{L}_{o}$ represents the Hamiltionian in our paper, and it acts on the Krylov basis and then we can obtain
$\mathcal{L}_{o}|\mathcal{O}_{n})=-ic_{n}|\mathcal{O}_{n})+b_{n+1}|\mathcal{O}_{n+1})+b_{n}|\mathcal{O}_{n-1}),$ (40)
where $c_{n}$ encodes the information of the open system which is the diagonal part and $b_{n}$ is the Lanczos coefficient and Eq. (40) was first defined by Bhattacharya:2023zqt . The expression of Lindbladian $\mathcal{L}_{o}$ in open system is as follows
$\mathcal{L}_{o}=\mathcal{H}_{o}=\mathcal{H}_{close}+\mathcal{H}_{open}.$ (41)
In our paper, the Hamilton (Lindbladian) is shown in Eq. (30). In light of Li:2024kfm , the Hamiltonian can be divided into two parts: the open part (the part of open system)
${\mathcal{H}_{open}}=(\frac{a^{2}}{2k}V_{,\phi\phi}+k)(\hat{c}_{anc}^{\dagger}\hat{c}_{anc}+\hat{c}_{\vec{k}}\hat{c}^{\dagger}_{\vec{k}})$ (42)
and the closed part (the part of closed system)
$\mathcal{H}_{close}=(\frac{a^{2}}{2k}V_{,\phi\phi}+i\frac{{a}^{\prime}}{a})\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{anc}^{\dagger}+(\frac{a^{2}}{2k}V_{,\phi\phi}-i\frac{{a}^{\prime}}{a})\hat{c}_{\vec{k}}\hat{c}_{anc}$ (43)
where $|\mathcal{O}_{n})=\left|n\right\rangle\otimes\left|n\right\rangle_{anc}$. In order to test accuracy of decomposition of Hamiltonian, one can use the open part of the Hamiltonian to plug into Eq. (40) since the target quantum state is $|\mathcal{O}_{n})$, one can straightforwardly obtain $c_{n}=i(2n+1)(\frac{a^{2}}{2k}V_{,\phi\phi}+k)$ and meanwhile the target quantum state is still $|\mathcal{O}_{n})$ after acting the Hamilton (40) towards $|\mathcal{O}_{n})$. Here, we just list the resulting formulas of $c_{n}$ and $b_{n}$ as following:
$c_{n}=i(2n+1)(\frac{a^{2}}{2k}V_{,\phi\phi}+k),\left|b_{n}\right|=n\sqrt{(\frac{a^{2}}{2k}V_{,\phi\phi})^{2}+(\frac{{a}^{\prime}}{a})^{2}}.$ (44)
The two quantities $c_{n}$ and $b_{n}$ in Eq. (44) are determined by Hamiltonian which are independent of $T_{k}$, and $c_{n}$ encodes the information of dissipative coefficient, $b_{n}$ measures the operator growth of one system. Also following Bhattacharya:2022gbz , we can find
$b^{2}_{n}=|1-\mu^{2}_{1}|n(n-1+\beta),c_{n}=i\mu_{2}(2n+\beta),$ (45)
where we already defined two parameters $\mu_{1}$ and $\mu_{2}$ ($\mu_{1}=\mu_{2}$ corresponds to $\mu$ of Bhattacharya:2022gbz ) and $\mu_{2}$ plays a role of dissipative coefficient. Combining Eqs. (44) and (45), one could find the following correspondence,
$\left|1-\mu_{1}\right|^{2}=(\frac{a^{2}}{2k}V_{,\phi\phi})^{2}+(\frac{{a}^{\prime}}{a})^{2},\mu_{2}=(\frac{a^{2}}{2k}V_{,\phi\phi}+k).$ (46)
where $\beta=1$ in our case and one can observe that $\mu_{1}$ and $\mu_{2}$ are exactly determined by Hamiltonian (30) which is not like Ref Bhattacharya:2022gbz . Here, we will give a little more about $b_{n}$. Ref. Parker:2018yvk has proposed that the Lanczos coefficient will be satisfied with the following relation in the thermal limit,
$b_{n}\leq\alpha n+\eta,$ (47)
where $\alpha$ and $\eta$ encodes the information of many-body system, when $b_{n}=\alpha n+\eta$ that means the system will reach the maximal chaos, we could find that $\eta=0$ in our work and $\alpha$ is encoded in Eq. (44) whose formula is $\sqrt{(\frac{a^{2}}{2k}V_{,\phi\phi})^{2}+(\frac{{a}^{\prime}}{a})^{2}}$. Meanwhile we could find that the whole early universe is a perfect chaotic system since $b_{n}\propto n$ leads to the maximal chaos via Bhattacharya:2022gbz , furthermore the Lyapunov index $\lambda$ could relate to $\alpha$ as follows,
$\lambda=2\alpha,$ (48)
where there is a simple relation $b_{n}=2\alpha=\lambda$ as $n=2$, the Lanczos coefficient represents the chaotic feature of the system. Consequently, we only plot $b_{n}$ as a manifestation for illustrating the chaos and operator growth.
Figure 5: The numeric of $b_{n}$ in terms of $\eta$ for three different periods (inflation, $RD$ and $MD$), in which we have set $b_{n}=2\alpha=\lambda$ as $n=2$, $H_{0}=10^{-2.5}$ and $m_{eff}=10^{-6}$ for simplicity. The momentum can be set various values in distinctive periods.
Fig. 5 shows the evolution of $b_{n}$ varying concerning conformal time $\eta$ for the entire early universe, where momentum decreases from inflation to MD. For the first panel, we could see that the $b_{n}$ is enhancing which means that the chaos is always increasing due to its definition. Meanwhile, the formula of Lanczos coefficient is $b_{n}=n\frac{a^{\prime}}{a}$ independent of $k$ where we have neglected the contribution of potential due to the slow-roll conditions. Once entering the RD and MD, we could see that the total trend of chaos will be decreased during these two periods, where the chaos will be a little enhanced as enhancing $k$ but they were still in the same energy scale. What reasons lead to this kind of phenomenon, one reason is that the contribution of inflationary potential is taken into account due to the violation of slow-roll conditions, and the second reason is that the evolution of the scale factor is different compared with inflation. Recalling the preheating process (in RD and MD), the generation of particles will decrease the chaos in MD and then the chaos will be enhanced a little bit by improving the momentum where it occurs at some certain scales.
7.2 The wave function with open system’s method
We will follow the method of Parker:2018yvk for constructing the wave function with the method of open system. Before the procedure, we will discuss its validity that is based on the asymptotic behavior of $b_{n}^{(1)}$
$b_{n}^{(1)}=\alpha n+\gamma,$ (49)
where it represents the maximal chaos that is consistent with calculation as shown in (44), which is valid for the infinite, non-integrable, many-body chaotic system, compared with the original method of Parker:2018yvk , we have defined two parameters $\mu_{1}$ and $\mu_{2}$, showing in Eq. (45), which are exactly determined by Hamiltonian (46). As $\mu_{1}=\mu_{2}=\mu$, it will nicely recover the traditional method. In Ref. Parker:2018yvk , they focused on one parameter $\mu$ where they only consider some constant values ($\mu\ll 1$ meaning the weak dissipative system). Through Eq. (45), the advantage of our method is that we could proceed with more realistic models including various quantum inflationary models, we only concentrate on the single field inflation in this work. Based on these considerations, Ref. Li:2024kfm has constructed the wave function by extending the original method, which is denoted by,
$\phi_{n}=\frac{\operatorname{sech}\eta}{1+\mu_{2}\tanh\eta}|1-\mu^{2}_{1}|^{\frac{n}{2}}(\frac{\tanh\eta}{1+\mu_{2}\tanh\eta})^{n},$ (50)
where our previous study Li:2024kfm has tested its validity by calculating Kryov complexity and and K-entropy under the weak dissipative approximation, which could nicely recover into the case with the method of closed system. During this process of calculation, there is correspondence between the leading order of Krylov complexity within the method of open system and the Krylov complexity within the approach of closed system, which $\eta$ is corresponding to $r_{k}$ and thus we will keep the temperature in the following the construction of wave function taking place of $\eta$. Being armed with this correspondence, and meanwhile we make use of $\beta\omega=-\ln\tanh^{2}r_{k}$, we can similarly construct the wave function as follows within the method of open system,
$\phi_{n}=\frac{\sqrt{1-e^{-\frac{\omega}{k_{B}T}}}}{1+\mu_{2}e^{-\frac{\omega}{2k_{B}T}}}|1-\mu^{2}_{1}|^{\frac{n}{2}}(\frac{e^{-\frac{\omega}{2k_{B}T}}}{1+\mu_{2}e^{-\frac{\omega}{2k_{B}T}}})^{n}$ (51)
where $\operatorname{sech}\eta\sim\operatorname{sech}{r}_{k}=\sqrt{1-e^{-\frac{\omega}{k_{B}T}}}$, $\tanh\eta\sim\tanh{r_{k}}=e^{-\frac{\omega}{2k_{B}T}}$. Subsequently, we could calculate the Krylov complexity in light of (10) whose resulting formula is
$K=\frac{|1-\mu_{1}^{2}|e^{-\frac{\omega}{k_{B}T}}(1-e^{-\frac{\omega}{k_{B}T}})}{[1+2\mu_{2}e^{-\frac{\omega}{2k_{B}T}}+(\mu_{2}^{2}-|1-\mu_{1}^{2}|)e^{-\frac{\omega}{k_{B}T}}]^{2}}.$ (52)
Being armed with the correspondence $\beta\omega=-\ln\tanh^{2}r_{k}$, we could check the Krylov complexity of thermal state is consistent with Li:2024kfm . Within the weak dissipative approximation, namely $\mu_{1}\ll 1$ and $\mu_{2}\ll 1$, we could obtain the leading order of Eq. (52) as follows
$K=\frac{1}{2}e^{-\frac{\omega}{2k_{B}T}}\sinh^{-1}{(\frac{\omega}{2k_{B}T})}+\mathcal{O}^{n}(\mu_{2}),$ (53)
where the leading order of Eq. (53) is consistent with Krylov complexity (36) and $\mathcal{O}^{n}(\mu_{2})$ denotes the higher order term of Krylov complexity (52) in terms of $\mu_{2}$. Once obtaining the complete formula of Krylov complexity, we could numerically simulate its evolution in terms of conformal time $\eta$.
Figure 6: The numeric of Krylov complexity (52) in terms of $\eta$ for three different periods (inflation, $RD$ and $MD$), in which we have set $k_{B}=1$, $\omega=1$, $H_{0}=10^{-2.5}$ and $m_{eff}=10^{-6}$ for simplicity. The momentum and temperature can be set various values in distinctive periods.
Fig. 6 indicates the evolution of Krylov complexity (52) in three epochs, including inflation, RD, and MD. We have set different temperatures $T$ and $k$ in these three periods according to their corresponding energy scale. Let us analyze the panel of Fig. 6 individually. $(a)$. In inflation, one can see that the Krylov complexity is growing, while the Krylov complexity will grow slower as enhancing the temperature. $(b)$. Once entering RD, the Krylov complexity will decrease and then saturate to some constant values that are consistent with the investigations within the method of a closed system as shown in Fig. 3. Our numerical decreasing trend is mainly determined by temperature not by momentum $k$. Meanwhile, the evolution trend of Krylov complexity is similar to Fig. 3, the reason comes via the contribution of inflaton’s potential. $(c)$. As for MD, we could see that Krylov complexity is similar to RD, the only difference is that the Krylov complexity will be enhanced in some certain scales for the lower temperature, where we guess that the particle generation leads to this kind of evolution behavior of Krylov complexity.
7.3 K-entropy with the approach of open system
We already calculated the Krylov complexity in terms of (51), meanwhile we also checked its leading order under the weak dissipative approximation could nicely recover into the case of closed system’s method. For completeness, we will calculate the K-entropy by means of (51) and investigate the evolution of K-entropy. By the definition of K-entropy $S_{K}=-\sum^{\infty}_{n=0}|\phi_{n}|^{2}\ln|\phi_{n}|^{2}$, we could obtain the resulting formula as follows,
$\begin{split}S_{K}=&-(1-e^{-\frac{\omega}{k_{B}T}})\frac{(1+\mu_{2}e^{-\frac{\omega}{2k_{B}T}})^{2}[\ln{(1-e^{-\frac{\omega}{2k_{B}T}})}-2\ln{(1+\mu_{2}e^{-\frac{\omega}{2k_{B}T}})}]}{[1+2\mu_{2}e^{-\frac{\omega}{2k_{B}T}}+(\mu_{2}^{2}-|1-\mu_{1}^{2}|)e^{-\frac{\omega}{k_{B}T}}]^{2}}\\ &-(1-e^{-\frac{\omega}{k_{B}T}})\frac{|1-\mu_{1}^{2}|e^{-\frac{\omega}{k_{B}T}}[\ln{|1-\mu_{1}^{2}|}+e^{-\frac{\omega}{k_{B}T}}-\ln{(1-e^{-\frac{\omega}{k_{B}T}})}]}{[1+2\mu_{2}e^{-\frac{\omega}{2k_{B}T}}+(\mu_{2}^{2}-|1-\mu_{1}^{2}|)e^{-\frac{\omega}{k_{B}T}}]^{2}},\end{split}$ (54)
when $\mu_{1}\ll 1$ and $\mu_{2}\ll 1$, we could obtain the K-entropy of closed system as follows
$\begin{split}S_{K}=\frac{\omega}{2k_{B}T}e^{-\frac{\omega}{2k_{B}T}}\sinh^{-1}{(\frac{\omega}{2k_{B}T})}-\ln{[2\sinh{(\frac{\omega}{2k_{B}T}})]}+\frac{\omega}{2k_{B}T}+\mathcal{O}^{n}(\mu_{2}).\end{split}$ (55)
where it is consistent with Eq. (38). Based on these two checking for Krylov complexity and K-entropy, we could see that the construction of wave function (51) is correct.
Figure 7: The numeric of K-entropy (54) in terms of $\eta$ for three different periods (inflation, $RD$ and $MD$), in which we have set $k_{B}=1$, $\omega=1$, $H_{0}=10^{-2.5}$ and $m_{eff}=10^{-6}$ for simplicity. The momentum and temperature can be set various values in distinctive periods.
Being armed with K-entropy (54), Fig. 7 clearly indicates the complete evolution of K-entropy in the whole early universe, the parameters what we set are the same with Fig. 6. Keeping in mind of the meaning of entropy, we could see that the disorder will increase in inflation, especially for the end of inflation. Then, the chaos will decrease to some constant values. When the particles are generated via preheating, the chaos will saturate some other constant values. We could see that the evolution trend of K-entropy is similar to the K-entropy within the method of closed system as shown in 4, the difference come via that we already maintain the temperature explicitly.
7.4 Some discussions for $\mu_{2}$
In our previous study Li:2024kfm , we already concluded that the inflation is a strong dissipative system leading to $\mu_{2}\gg 1$, and we also guessed that the Krylov complexity will saturate some constant values in RD and MD which the RD and MD may be the weak dissipative region. In this section, we will confirm this point. First, from the evolution trend of Krylov complexity in both methods, the method of closed system and the method of open system. From Figs. 4 and 6, it indicates that the Krylov complexity will saturate some constant values which is consistent with the discussions of Ref. Parker:2018yvk in the weak dissipative region. Another reason is from the explicit formula of $\mu_{2}=(\frac{a^{2}}{2k}V_{,\phi\phi}+k)$, if we take values of parameters that are given in Figs. 4 and 6, we could see that $\mu_{2}\ll 1$ in RD and MD. Consequently, our numeric has confirmed our previous guess which the RD and MD are the weak dissipative system.
Let us make some comments for Sec. 7. In this section, we already used the method of open system to investigate the evolution of Krylov complexity in the entire early universe. Through the numeric, one can see the complete evolution of Krylov complexity and K-entropy. In inflation, the Krylov complexity will continue growing due to the strong dissipative nature of inflation. Once entering RD and MD, we could see that Krylov complexity will saturate some constant values since the RD and MD are the weak dissipative systems, where the key reason leading to this kind of behavior is that the contribution of potential for inflaton has been taken into account. As for the K-entropy, we could see that the chaos will be enhanced in inflation, and then the chaos will approach some constant values in RD and MD due to the generation of particles via preheating. Thus, we could conclude that the generation of particles will change the environment from the strong dissipative region to the weak dissipative region.
8 Summary and outlook
With the development of Krylov complexity, it becomes more significant in high-energy physics. The whole early universe includes inflation, RD, and MD which gives a platform for generating all of the matter including observed matter, dark matter, and dark energy. Thus, it is of particular interest to explore the Krylov complexity in the entire early universe. From another aspect, the Krylov complexity is free of ambiguities compared with circuit complexity. Therefore, we investigate the Krylov complexity due to its uniqueness in the definition. Meanwhile, in the whole early universe in a highly non-equilibrium system, the Krylov complexity of the thermal state will be more practical and reliable compared with the two-mode squeezed state. Once entering RD and MD, the contribution of inflationary potential cannot be ignored due to the violation of slow-roll conditions, where the potential is neglected by the slow-roll conditions. How the Krylov complexity varies in RD and MD is also essential, since it helps us construct a complete picture of the evolution of Krylov complexity in the whole early universe.
In this work, we have systematically investigated the Krylov complexity in both methods, including the method of closed system and the approach of open system, in the entire early universe. Here, we summarize our main results.
$(a)$. In Sec. 3, we give some fundamental knowledge about the entire early universe (inflation, RD, and MD) characterized by the scale factor as shown in Eq. (16). Since the single-mode thermal state is a mixed state whose wave function cannot be written down explicitly. Then, we define the wave function of the thermal state as the purification of the single thermal state as shown in Eq. (19) which can be dubbed as the TFD. Meanwhile, we could find the correspondence between the factor $r_{k}$ and temperature $T_{k}$. Consequently, our wave function with the thermal state can be considered as the two-mode pure state as shown in Eq. (21). In this work, we aim to investigate the impact of inflationary potential on the Krylov complexity. Therefore, we did not adopt the Mukhanov-Sasaki variable since it contains the contribution of inflationary potential. In Sec. 5, we just perturb the inflaton field and then make the covariant quantization for action (24). The final Hamilton can be written by (30), including the inflationary potential.
$(b)$. Following the method of closed system, we obtain the numeric of $T_{k}$ and $\phi_{k}$ in terms of conformal time $\eta$ as shown in Figs. 1 and 2. The most essential evolution is for the temperature since the Krylov complexity and K-entropy only depend on the $T_{k}$. Our numeric indicates that the temperature will be enhanced in inflation. In RD and MD, the temperature will be oscillating around some constant values. In light of Figs. 1 and 2, the numeric of Krylov complexity is shown in Fig. 3. The evolution trend of Krylov complexity is almost the same as the temperature. Physically speaking, the particles will be generated via preheating in RD and MD, whose energy comes via the oscillation around the minimum of the inflationary potential, the potential can be approximated to be the quadratic form as shown in Eq. (25). Once taking the contribution of inflationary potential into account, the Krylov complexity will saturate some constant values. As for the K-entropy, it follows a similar trend as shown in Fig. 4. Recalling that the entropy represents the chaos of the dynamical system, the chaos is enhanced in inflation, and then the chaos will approach some constant values in RD and MD, where the generation of particles leads to this kind of evolution.
$(c)$. In both methods (the method of closed system and the method of open system), there is an essential quantity called Lanczos coefficient representing the chaotic feature of one dynamical system. In our paper, it is exactly determined by the Hamiltonian as shown in Eq. (44). Since it also represents the chaos of one dynamical system, thus we can see that its trend is similar to K-entropy as shown in Fig. 5. Within the method of open system, we utilize the generalized method for constructing the thermal state as shown in Eq. (51) which is used to calculate the Krylov complexity (52) and K-entropy (54), where their leading order is consistent with the case of closed system’s method. Therefore, the construction of wave function (44) is correct. Being armed with the Krylov complexity (52) and K-entropy (54), we could obtain their numeric as shown in Figs. 6 and 7. Through these two figures, we can see that the trend of Krylov complexity and K-entropy is consistent with each other. In inflation, they continue growing and then they will approach some constant values in RD and MD. Meanwhile, we found that the RD and MD are weak dissipative systems. Therefore, we could conclude that the transition from the strong dissipative system to the weak dissipative system will lead to the Krylov complexity and K-entropy approaching some constant values, which is consistent with our previous discussions via Li:2024kfm .
Through the summary, we could see that a complete picture of the early universe. In inflation, the Krylov complexity will continue growing due to its strong dissipative nature. Once entering the RD and MD, they will become the weak dissipative system which the corresponding Krylov complexity will approach to some constant values. This kind of trend is also applied to their corresponding K-entropy. Although we already obtained the complete evolution of Krylov complexity in the whole universe. There are still lots of ideas worth investigating. Here, we just list some problems for the future research.
$(a)$. The Refs. Kofman:1997yn ; Kofman:1994rk ; Cheung:2007st has claimed that the universe is an open system. And the traditional method of dealing with the open system is using the non-Hermitian Hamiltonian. However, our Hamiltonian (30) is Hermitian, whereas our wave function (51) is not unitary anymore. To obtain the non-Hermitian Hamiltonian, one possibility is expanding the action (23) into a higher order, such as the third order in terms of the Mukhanov-Sasaki variable. Meanwhile, we could introduce a more complicated form of potential compared with our approximated quadratic potential.
$(b)$. During inflation, the curvature perturbation will transit from the quantum level to the classical level. Thus, we could consider how the decoherence Burgess:2022nwu impacts the Krylov complexity in RD and MD. In our paper, we only consider the single field inflation, we could extend our investigations into the muli-field inflation Liu:2019xhn ; Liu:2020zzv ; Liu:2021rgq ; Zhang:2022bde and $f(R)$ gravity Liu:2020zlr ; Liu:2018htf ; Liu:2018hno , which could test the geometrical effects of field space and non-trivial part of $f(R)$ for the Krylov complexity and K-entropy in RD and MD.
Acknowledgements.
LH and TL are supported by NSFC grant NO. 12165009, Hunan Natural Science Foundation NO. 2023JJ30487 and NO. 2022JJ40340.

### v2

@@
 The very recent developments of Krylov complexity can be found in Refs.
-Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj .
+Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj ; Camargo:2023eev ; Camargo:2024deu ; Baggioli:2024wbz .
 Especially, the detailed implication of Krylov complexity in various fields can be found in Ref.
@@
 4 The wave function of thermal state
-Physically speaking, the Hamiltonian of single-field inflation can be considered as an inverted harmonic oscillator Li:2024kfm .
-Meanwhile, we will investigate the Krylov complexity for the thermal state.
+We will investigate the Krylov complexity for the thermal state.
 $\begin{split}\hat{\rho}=\frac{1}{\sqrt{z}}\sum^{\infty}_{n=0}e^{-\beta E_{n}}|n\rangle\langle n|\end{split}$ (17)
 In the context of bosonic oscillators, where $\beta=\frac{1}{k_{B}T}$ and $E_{n}=n\omega$ ($\omega$ is the frequency), we will consider the thermal state (17) as a pure state following the approach in Jefferson:2017sdb ; Ali:2018fcz , using the method described in Haque:2021kdm .
-Given that the thermal state is a mixed state, we need to create the purification of $\hat{\rho}$, which we refer to as a pure state $|\Psi\rangle$, in an enlarged Hilbert space $\mathcal{H}_{pure}=\mathcal{H}\otimes\mathcal{H}_{anc}$, where $\mathcal{H}_{anc}$ represents the ”ancillary” degrees of freedom.
+Since the thermal state is a mixed state, we need to create the purification of $\hat{\rho}$, which refers as a pure state $|\Psi\rangle$, in an enlarged Hilbert space $\mathcal{H}_{pure}=\mathcal{H}\otimes\mathcal{H}_{anc}$, where $\mathcal{H}_{anc}$ represents the ”ancillary” degrees of freedom.
 When we trace $|\Psi\rangle$ over $\mathcal{H}_{anc}$, we obtain the mixed state $\hat{\rho}_{mix}$, indicating that $|\Psi\rangle$ serves as the purification of the thermal state (17).
+Similar method can also be found in Das:2024zuu , which they can also map the pure state into the mixed state.
 It is important to note that the observable is preserved even in the purification process, as evidenced by $\hat{\mathcal{O}}=\rm Tr_{anc}\langle\Psi|\mathcal{O}|\Psi\rangle=\rm Tr\big{(}\hat{\rho}_{mix}\hat{\mathcal{O}}\big{)}$, which maintains the reliability of calculating the Krylov complexity using this approach.
@@
 […]
-With the application of Fourier decomposition,
+With the application of Fourier decomposition and Fourier transformation,
 $\hat{v}(\eta,\vec{x})=\int\frac{d^{3}k}{(2\pi)^{3/2}}\sqrt{\frac{1}{2k}}(\hat{c}_{\vec{k}}^{\dagger}v_{\vec{k}}^{\ast}(\eta)e^{-i\vec{k\cdot}\vec{x}}+{\hat{c}_{\vec{k}}}v_{\vec{k}}e^{i\vec{k\cdot}\vec{x}}),$ (28)
 $\hat{\pi}(\eta,\vec{x})=i\int\frac{d^{3}k}{(2\pi)^{3/2}}\sqrt{\frac{k}{2}}(\hat{c}_{\vec{k}}^{\dagger}u_{\vec{k}}^{\ast}(\eta)e^{-i\vec{k\cdot}\vec{x}}-\hat{c}_{\vec{k}}u_{\vec{k}}e^{i\vec{k\cdot}\vec{x}}).$ (29)
-As a result, we could obtain the Hamiltonian in terms of creation and annihilation operators,
+As a result, we could obtain the Hamiltonian in terms of creation and annihilation operators in momentum space,
 $\begin{split}\hat{H}_{k}&=(\frac{a^{2}}{2k}V_{,\phi\phi}+k)(\hat{c}_{anc}^{\dagger}\hat{c}_{anc}+\hat{c}_{\vec{k}}\hat{c}^{\dagger}_{\vec{k}})+(\frac{a^{2}}{2k}V_{,\phi\phi}+i\frac{{a}^{\prime}}{a})\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{anc}^{\dagger}\\ &+(\frac{a^{2}}{2k}V_{,\phi\phi}-i\frac{{a}^{\prime}}{a})\hat{c}_{\vec{k}}\hat{c}_{anc}\end{split}$ (30)
@@
 4 and 6, we could see that $\mu_{2}\ll 1$ in RD and MD.
+Physically speaking, the energy of inflationary potential will transfer into the other field that is dubbed as preheating, where this process will lead to the generation of particles, we could conclude that the process of preheating in MD will make the universe transit from the strong dissipative region to the weak dissipative region.
 Consequently, our numeric has confirmed our previous guess which the RD and MD are the weak dissipative system.